OpenStudy (anonymous):

Help with one slope points problem? Find the equation of the line below. Arrange your answer in the form y = mx + b, where b is the constant. Points (-2,2) and (2,-2) and my answer is y = x.. it doesn't seem right.

OpenStudy (anonymous):

i dont think its right there

OpenStudy (campbell_st):

do you know the formula for finding the slope of a line given 2 points...?

OpenStudy (anonymous):

Find the slope of the line...

OpenStudy (anonymous):

Yes. I do know the formula...

OpenStudy (anonymous):

The slope is 4, right?

OpenStudy (anonymous):

Then substitute the values in the formula and find the slope first...

OpenStudy (anonymous):

No it is not right...

OpenStudy (campbell_st):

well looking at your solution, the gradient is correct... so the line is y = x + b to find the value of b, just substitute the x and y values from either point and solve for b

OpenStudy (anonymous):

Can you show your work so that we can find where you went wrong ??

OpenStudy (campbell_st):

oops the slope is -1 not 1 so the line is y = -x + b

OpenStudy (anonymous):

Okay.. 2 = -2 + b?

OpenStudy (anonymous):

What next?

OpenStudy (anonymous):

What are you doing ???

OpenStudy (anonymous):

Have you find the slope ???

OpenStudy (anonymous):

*found.

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

What is slope according to the formula you know ??

OpenStudy (anonymous):

\[Slope = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

That's where I'm stuck... I thought it was 4 down and 4 across, wait, is the slope 0?

OpenStudy (anonymous):

No it's not..

OpenStudy (anonymous):

Just substitute the values in the above slope formula and be careful with negative signs..

OpenStudy (anonymous):

So I have -4 and positive 4....

OpenStudy (anonymous):

slope = 8?

OpenStudy (anonymous):

Yes.. What you get on dividing them ??

OpenStudy (anonymous):

um. what do I divide them by?

OpenStudy (anonymous):

\[y = \frac{-2 - (2)}{2 - (-2)} \implies \frac{-4}{4} \implies ??\]

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

Then what is the problem you are facing in finding the slope. This is the standard procedure for finding the slope.. Getting??

OpenStudy (anonymous):

Understood.

OpenStudy (anonymous):

Just remember the formula It will help you in lot of problem like we are doing..

OpenStudy (anonymous):

Now: y = mx + b Or by putting m = -1 (m is notation for Slope) y = -x + b Is it okay ??

OpenStudy (anonymous):

ok, y = -1x + b or y = -x + b

OpenStudy (anonymous):

Yep.. Now : Substitute any of two point given in it to find b...

OpenStudy (anonymous):

so I can use -2 and 2?

OpenStudy (anonymous):

Your first point is (-2,2).. Here x = -2 and y = 2..

OpenStudy (anonymous):

If you don't want to use the first point then use the second simple..!! (2, -2) here x = 2 and y = -2.. It is your wish to choose any one you want...

OpenStudy (anonymous):

Now substitute any one point and find b... You will get b same for both the cases...

OpenStudy (anonymous):

so if 2 = -2 + b.. is b a 4?

OpenStudy (anonymous):

If you are not getting then tell me, I will follow another method to make your understand this..

OpenStudy (anonymous):

I honestly don't understand.

OpenStudy (anonymous):

What is the step you are not getting ??

OpenStudy (anonymous):

Have you got how we found slope ??

OpenStudy (anonymous):

I understand how to find the slope, but this confused me: Now substitute any one point and find b... You will get b same for both the cases...

OpenStudy (anonymous):

Ha ha ha ha.. Okay.. I just said if you put first point or if you put second point, you will get the same value for b in both the cases..

OpenStudy (anonymous):

Okay, that helps a bit.. so what points do I use, and after that, how do I find the value for b?

OpenStudy (anonymous):

Just forget what I just said.. Tell me what is first point given to you ??

OpenStudy (anonymous):

Um, let's see.. .(-2,2)

OpenStudy (anonymous):

Yep.. Now tell me what is x and what is y here ??

OpenStudy (anonymous):

x is -2, y is 2.

OpenStudy (anonymous):

Now you have ater putting slope : y = -x + b Right ?? Can you substitute here x and y values to find b ??

OpenStudy (anonymous):

*after

OpenStudy (anonymous):

2 = -2 + b

OpenStudy (anonymous):

ok, then..

OpenStudy (anonymous):

Yep.. Now 2 both the sides and find b here..

OpenStudy (anonymous):

b = 4

OpenStudy (anonymous):

Yep.. Now you have again : y = mx + b Right ??

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

Here, you have found m = -1 and b = 4 (b you have just found).. Here put these values in the equation : y = mx + b and tell me what do you get ?? Don't put values of x and y here.. Just leave them as x and y.. Tell me what did you get??

OpenStudy (anonymous):

y = -1x + b

OpenStudy (anonymous):

or y = -x + b

OpenStudy (anonymous):

Substitute for b also..

OpenStudy (anonymous):

y = -x + 4

OpenStudy (anonymous):

You have found the value of b just above you can safely put it here.. Yep.. So: y = -x + 4 is your required equation..

OpenStudy (anonymous):

well, thanks so much for your help.

OpenStudy (anonymous):

Welcome dear..